Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used iRequired: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overhRequired: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overhRequired: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overhRequired: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overh
Answer #1
Required 1 :
Predetermined overhead rate = Estimated manufacturing overhead / Estimated direct materials costs = 800000 / 500000 160%
Required 2 :
Beginning raw materials 20000
(+) Purchase of raw materials 510000
Total raw materials available 530000
(-) Ending raw materials 80000
Raw materials used 450000
Overhead applied = Raw materials used * Predetermined overhead rate = 450000 * 160% 720000
Indirect labor 170000
Property taxes 48000
Depreciation of equipment 260000
Maintenance 95000
Insurance 7000
Rent, building 180000
Actual overhead costs 760000
As the overhead applied are less than the actual overhead costs, the overheads are underapplied
Underapplied overhead = Actual overhead costs - Overhead applied = 760000 - 720000 40000
Underapplied Overhead 40000
Required 3 :
Gitano Products
Schedule of cost of goods manufactured
Direct materials :
Beginning raw materials 20000
Purchase of raw materials 510000
   Total raw materials available 530000
Ending raw materials 80000
Raw materials used in production 450000
Direct labor cost 90000
Manufacturing overhead applied 720000
Total manufacturing costs 1260000
Beginning work in process 150000
Total cost of work in process 1410000
Ending work in process 70000
Cost of goods manufactured 1340000
Required 4 :
Cost of goods manufactured 1340000
(+) Beginning finished goods 260000
(-) Ending finished goods 400000
Unadjusted cost of goods sold 1200000
Required 5 :
Direct materials 24000
Direct labor cost ( 70000 - 24000 - 38400 ) 7600
Manufacturing overhead ( 24000 * 160% ) 38400
Work in process inventory 70000